OpenStudy (anonymous):

What is the solution to the equation (1 /√8) = 4^(m – 3) ? What is the solution to the equation (1 /√8) = 4^(m – 3) ? @Mathematics

OpenStudy (anonymous):

m = 5 / 4 m = –11 / 4 m = 9 / 4 m = –15 / 4

OpenStudy (anonymous):

2^(-3/2) =2^2(m-3) -3/2 =2m-6 etc

OpenStudy (anonymous):

prob -15/4

OpenStudy (anonymous):

Ok then if its like that -3/2 =2m-6 there would be 2 fraction bars.

OpenStudy (anonymous):

no not - 15.4 2m= 9/2 m= 9/4

OpenStudy (anonymous):

How did you get that?